### v2

# When do pseudo-Gorenstein rings become Gorenstein?

Abstract.
For a Cohen–Macaulay positively graded ring $R$, we say that $R$ is pseudo-Gorenstein if its leading coefficient is 1. In this paper, we study the relationship between canonical trace and pseudo-Gorensteinness for a graded ring. In particular, we show that if a nearly Gorenstein graded domain satisfies certain mild assumptions and is pseudo-Gorenstein, then it is necessarily Gorenstein. As an application, we clarify the relationships among nearly Gorensteinness, almost Gorensteinness, and levelness, which generalize the notion of Gorensteinness, in the context of standard graded domains. Moreover, we give a method for constructing quasi-Gorenstein rings by taking a Veronese subalgebra of certain Noetherian graded rings.
Key words and phrases:
Pseudo-Gorenstein, nearly Gorenstein, almost Gorenstein, level, semi-standard graded rings, trace ideals, quasi-Gorenstein, Veronese subalgebra
2020 Mathematics Subject Classification:
Primary 13H10, 13A02; Secondary 05E40.
1. Introduction
A Cohen–Macaulay standard graded ring $R$ over a field $\mathbb{k}$ is called pseudo-Gorenstein ([ene2015pseudo]) if the leading coefficient $h_{s(R)}$ of $R$ is 1. The structural analysis of pseudo-Gorensteinness within certain classes of rings has been actively studied (see, e.g., [hatasa2024pseudo, ene2015pseudo, rinaldo2024level, rinaldo2023level]). For simplicity, we will refer to a positively graded ring $R=\bigoplus_{i\geq 0}R_{i}$ simply as a graded ring. Now we discuss the pseudo-Gorenstein property. For a Cohen–Macaulay graded ring $R$, we say that $R$ is pseudo-Gorenstein if $h_{s(R)}=1$.
From Stanley’s classical result about $h$-vectors ([stanley1978hilbert, Theorem 4.1]), every Gorenstein graded ring is pseudo-Gorenstein. However, pseudo-Gorensteinness does not imply Gorensteinness in general. For example, all numerical semigroup rings are pseudo-Gorenstein, but there are infinitely many non-Gorenstein numerical semigroup rings. Is there a class of pseudo-Gorenstein rings which become Gorenstein?
Level rings, introduced by Stanley ([stanley2007combinatorics]), are among the earliest generalizations of Gorenstein rings. According to [stanley2007combinatorics, Chapter III, 3.2 Proposition], a Cohen–Macaulay graded ring $R$ is called level if the number of minimal generating system of the canonical module $\omega_{R}$ is $h_{s(R)}$. Therefore, every level and pseudo-Gorenstein ring is Gorenstein. As another generalization, Nearly Gorenstein rings were introduced by Herzog et al. ([herzog2019trace]). In this paper, we show that pseudo-Gorensteinness leads to Gorensteinness under assumptions weaker than nearly Gorensteinness. We refer the reader to Section missing 2 for notation.
Theorem A.
Let $R$ be a Cohen–Macaulay generically Gorenstein graded ring. Suppose that
$\text{grade}(R[\omega_{R}]_{-a_{R}})>0,\;\text{grade}(R[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}({\mathfrak{M}}_{R})})\geq 2.$
Then $R$ is pseudo-Gorenstein if and only if $R$ is Gorenstein.
The condition on the grade of the canonical trace appearing in Theorem A is a weakened version of the nearly Gorenstein property, and it is significantly less restrictive than the nearly Gorenstein property. We also provide a proof that applies to (not necessarily Cohen–Macaulay) Noetherian rings (Theorem 3.2).
Since the introduction of nearly Gorenstein rings, the structural analysis of these rings and the trace ideals of canonical modules has garnered significant attention, becoming an active area of research in recent years ([miyazaki2021Gorenstein, hall2023nearly, dao2020trace, miyashita2024linear, herzog2019trace, herzog2019measuring, celikbas2023traces, ficarra2024canonical, ficarra2024canonical!, caminata2021nearly, kumashiro2023nearly, bagherpoor2023trace, srivastava2023nearly, lu2024chain, lyle2024annihilators, miyazaki2024non, jafari2024nearly, moscariello2025nearly, miyashita2024canonical]). Nearly Gorenstein rings are Gorenstein on the punctured spectrum ([herzog2019trace, Proposition 2.3 (a)]), and thus, in dimensions at least one, they are generically Gorenstein. On the other hand, Theorem 3.2 demonstrates that under a condition far weaker than the nearly Gorenstein property—one that does not even require the ring to be generically Gorenstein—pseudo-Gorensteinness implies Gorensteinness. In particular, if a nearly Gorenstein domain is pseudo-Gorenstein, then under a certain assumption (if we remove it, there is a counterexample; see Section missing 3), it must be Gorenstein (Section missing 1). Theorem A and the consequences derived from it, as described above, were known in the case of semi-standard graded affine semigroup rings (see [miyashita2024comparing, Theorem 3.5, Corollary 3.6]).
Among the major generalizations of Gorensteinness is the concept of almost Gorensteinness ([goto2013almost, goto2015almost, barucci1997one]), levelness and nearly Gorensteinness. The relationships among these three properties have been actively investigated across various classes of rings ([herzog2019trace, miyashita2024comparing, moscariello2021nearly, miyashita2024canonical, hall2023nearly, miyashita2023nearly, higashitani2022levelness]). In the one-dimensional case, it has been established that every almost Gorenstein ring is necessarily nearly Gorenstein ([herzog2019trace, Proposition 6.1]). Furthermore, for a local ring with an infinite residue field and positive Krull dimension, having minimal multiplicity implies that being nearly Gorenstein also guarantees that the ring is almost Gorenstein ([herzog2019trace, Theorem 6.6]). However, in the case of higher-dimensional rings or those without minimal multiplicity, the relationships among these properties are still not well understood.
As mentioned later, level rings have a strong connection with pseudo-Gorenstein rings. Similarly, it is known that almost Gorenstein rings also have a strong relationship with pseudo-Gorenstein rings in the case of standard graded domains ([higashitani2016almost, Theorem 4.7]). Theorem A provides a new perspective on the relationship between nearly Gorenstein rings and pseudo-Gorenstein rings. By focusing on the classification based on the pseudo-Gorenstein property, we clarify how the three properties in question relate to one another in domains of higher dimension.
[…]
Let $R$ be a Cohen–Macaulay graded domain with a field $R_{0}=\mathbb{k}$. In (2) below, we assume that $\mathbb{k}$ is algebraically closed with $\operatorname{char}(\mathbb{k})=0$. Then the following holds:
  • (1)
    If $0pt(\mathbb{k}[R_{\text{indeg}({\mathfrak{M}}_{R})}])\geq 2$ (e.g., $R$ is a semi-standard graded ring with $\dim(R)\geq 2$) and
    $\sqrt{R[\textrm{tr}_{R}(\omega_{R})]_{{\text{indeg}({\mathfrak{M}}_{R})}}}\supseteq{\mathfrak{M}}_{R}$
    (e.g. $R$ is a nearly Gorenstein ring). Then $R$ is pseudo-Gorenstein if and only if $R$ is Gorenstein. In particular, if the Cohen–Macaulay type of $R$ is 2, then $R$ is level;
  • (2)
    Let $R$ be a nearly Gorenstein and almost Gorenstein standard graded domain over a field $\mathbb{k}$. If $\dim(R)\geq 2$ and $s(R)\geq 2$, then $R$ is Gorenstein. In particular, all nearly Gorenstein and almost Gorenstein standard graded domains over $\mathbb{k}$ of dimension $\geq 2$ are level.
The applications of the main result to Cohen–Macaulay rings have been described above. On the other hand, Theorem A can be formulated for non-Cohen–Macaulay rings (see Section missing 3). Finally, as one such application, we discuss how a quasi-Gorenstein ring can be constructed from a non-Cohen–Macaulay ring that satisfies analogous properties of nearly Gorensteinness and pseudo-Gorensteinness. If a standard graded ring $R$ of positive dimension is Gorenstein, any Veronese subalgebra of $R$ is nearly Gorenstein (see [herzog2019trace, Corollary 4.7]), however, we cannot construct a Gorenstein ring from the Veronese subalgebra of a non-Gorenstein Cohen–Macaulay standard graded ring ([goto1978graded, Theorem (3.2.1)]). In another direction, we show that by considering the analogous properties of pseudo-Gorenstein (resp. nearly Gorenstein) rings for a Noetherian ring, one can construct quasi-Gorenstein rings from certain Veronese subalgebras.
[…]
Let $R$ be a Noetherian semi-standard graded generically Gorenstein ring such that $[\omega_{R}]_{-a_{R}}$ contains an $R$-regular element. Suppose that $a_{R}\in k\mathbb{Z}$ and $0pt(R^{(k)})\geq 2$ for some $k\in\mathbb{Z}_{>0}$. Additionally, assume that $\sqrt{[\textrm{tr}_{R}(\omega_{R})]_{1}R}\supseteq{\mathfrak{M}}_{R}$ (e.g., $\textrm{tr}_{R}(\omega_{R})\supseteq{\mathfrak{M}}_{R}$) and $\dim_{\mathbb{k}}([\omega_{R}]_{-a_{R}})=1$. Then $R^{(k)}$ is quasi-Gorenstein. In particular, if $R^{(k)}$ is Cohen–Macaulay, then it is Gorenstein.
Outline
In Section missing 2, we prepare some foundational facts and definitions for later discussions, collecting basic definitions and properties about generalizations of the Gorenstein property. In Section missing 3, we prove the general statement Theorem 3.1 for $R$-modules that are not necessarily finitely generated over a graded ring $R$. As a consequence, we establish Theorem A. In Section missing 4, as an application of Theorem A, we discuss various corollaries derived from it. First, we clarify the relationships among nearly Gorensteinness, almost Gorensteinness and levelness in the setting of standard graded domains (Section missing 1). Furthermore, we explore rings that are non-Cohen–Macaulay but satisfy an analogue of the nearly Gorenstein property, demonstrating that their Veronese subalgebras yield quasi-Gorenstein rings (Section missing 1).
2. Preliminaries
Throughout this paper, we fix a field $\mathbb{k}$ and denote the set of integers by $\mathbb{Z}$. The purpose of this section is to lay the groundwork for the proof of our main results. In this paper, we will refer to a positively graded ring $R=\bigoplus_{i\geq 0}R_{i}$ simply as a graded ring. Moreover, for any graded ring $R$, we always assume that $R_{0}=\mathbb{k}$. Notice that every Noetherian graded ring $R=\bigoplus_{i\geq 0}R_{i}$ has a unique graded maximal ideal ${{\mathfrak{M}}_{R}}$ and admits a graded canonical module $\omega_{R}$. Let $a_{R}$ denote the $a$-invariant of $R$. Note that $a_{R}=-\operatorname{min}\{j:(\omega_{R})_{j}\neq 0\}$. For a $\mathbb{Z}$-graded $R$-module $M$ and $k\in\mathbb{Z}$, let $M(k)$ denote the graded $R$-module having the same underlying $R$-module structure as $M$, where $[M(k)]_{n}=[M]_{n+k}$ for all $n\in\mathbb{Z}$. Moreover, we set
$\text{Hom}^{i}_{R}(M,R):=\{\phi\in\text{Hom}_{R}(M,R):\phi(M_{k})\subseteq R_{k+i}\text{\;for any $k\in\mathbb{Z}$}\}.$
Let $\textrm{tr}_{R}(M)$ denote the sum of ideals $\phi(M)$ with $\phi\in\text{Hom}^{*}_{R}(M,R):=\bigoplus_{i\in\mathbb{Z}}\text{Hom}^{i}_{R}(M,R)$ (Note that $\text{Hom}^{*}_{R}(M,R)=\text{Hom}_{R}(M,R)$ if $M$ is finitely generated). Namely,
$\textrm{tr}_{R}(M)=\sum_{\phi\in\text{Hom}^{*}_{R}(M,R)}\phi(M).$
[…]
It is known that $\textrm{tr}_{R}(\omega_{R})$ describes the non-Gorenstein locus if $R$ is Cohen–Macaulay (see [herzog2019trace, Lemma 2.1] or [HerzogKunz, Lemma 6.19]). In particular, $R$ is Gorenstein if and only if $\textrm{tr}_{R}(\omega_{R})=R$. Let $R$ be a graded ring and $I$ a graded ideal of $R$ containing a non-zero divisor of $R$. Set $I^{-1}:=\{x\in Q(R):xI\subseteq R\}$, where $Q(R)$ is the total quotient ring of fractions of $R$. Then we have $\textrm{tr}(I)=I\cdot I^{-1}$ (see [herzog2019trace, Lemma 1.1]).
[…]
For a finitely generated $R$-module $M$ over a Noetherian graded ring $R$, a maximal $M$-regular sequence (see [bruns1998cohen, Definition 1.1.1]) in ${\mathfrak{M}}_{R}$ always exists, and moreover, its length is constant (see [bruns1998cohen, Theorem 1.2.5]). We denote this length by $\text{grade}(M)$. The depth of $R$, denoted by $0pt(R)$, is defined as the length of a maximal $R$-regular sequence in the $R$-module $R$. This length is also written as $\text{grade}(R)$.
[…]
Let $R=\bigoplus_{i\geq 0}R_{i}$ be a Noetherian graded ring. If $R=\mathbb{k}[R_{1}]$, that is, $R$ is generated by $R_{1}$ as a ring, then we say that $R$ is a standard graded ring. If $R$ is finitely generated as a $\mathbb{k}[R_{1}]$-module, then we say that $R$ is a semi-standard graded ring.
The Ehrhart rings of lattice polytopes and the face rings of simplicial posets (see [stanley1991f]) are typical classes of semi-standard graded rings. From the perspective of combinatorial commutative algebra, the concept of semi-standard graded rings naturally arises in this context.
Let $R$ be a Noetherian graded ring of Krull dimension $d$. Suppose that $R$ (possibly after extending the ground field $\mathbb{k}$, see [stanley1978hilbert, 1.1. Lemma (vii)]) has a homogeneous $R$-regular sequence whose elements have degrees $r_{1},\cdots,r_{d}$. Then the Hilbert series $H(R,t)$ of $R$ is of the form
$H(R,t):=\sum_{i\geq 0}(\dim_{\mathbb{k}}(R_{i}))t^{i}=\frac{h_{0}+h_{1}t+\cdots+h_{s}t^{s}}{(1-t^{r_{1}})\cdots(1-t^{r_{d}})}$
for some integers $h_{0},h_{1},\cdots,h_{s}$ with $\sum_{i=0}^{s}h_{i}\neq 0$ and $h_{s}\neq 0$. We call the integer sequence $(h_{0},h_{1},\cdots,h_{s})$ the $h$-vector of $R$ and denote it as $h(R)$. Moreover, we call $s$ the socle degree of $R$ and denote it as $s(R)$. We always have $h_{0}=1$ because $h_{0}=H(R,0)=\dim_{\mathbb{k}}(R_{0})$. Moreover, if $R$ is Cohen–Macaulay, its $h$-vector satisfies $h_{s(R)}=\dim_{\mathbb{k}}([\omega_{R}]_{-a_{R}})$ and $h_{i}\geqq 0$ for every $i$. For further information on the $h$-vectors of Noetherian graded rings, see [stanley1978hilbert] and [stanley2007combinatorics]. We collects the definitions of generalizations of Gorenstein graded rings below.
[…]
Let $R$ be a Cohen–Macaulay graded ring.
  • (1)
    We say that $R$ is pseudo-Gorenstein ([ene2015pseudo]) if $h_{s(R)}=1$. Since $h_{s(R)}=\dim_{\mathbb{k}}([\omega_{R}]_{-a_{R}})$, This can also be rephrased as $\dim_{\mathbb{k}}([\omega_{R}]_{-a_{R}})=1$. While [ene2015pseudo] defines this term for Cohen–Macaulay standard graded rings, we use the term for (Cohen–Macaulay) graded rings.
  • (2)
    $R$ is called level ([stanley2007combinatorics]) if $\omega_{R}$ is generated in a single degree. This is equivalent to the number of minimal generators of $\omega_{R}$ being exactly $h_{s(R)}$ ([stanley2007combinatorics, Chapter III, 3.2 Proposition]).
  • (3)
    $R$ is called nearly Gorenstein ([herzog2019trace, Definition 2.2]) if $\textrm{tr}(\omega_{R})\supseteq{{\mathfrak{M}}_{R}}$. In particular, $R$ is nearly Gorenstein but not Gorenstein if and only if $\textrm{tr}(\omega_{R})={\mathfrak{M}}_{R}$. Moreover, based on [miyashita2024linear], we say that $R$ satisfies ()($\natural$)( ♮ ) if $R$ is semi-standard graded and $\sqrt{R[\textrm{tr}_{R}(\omega_{R})]_{1}}\supseteq{\mathfrak{M}}_{R}$.
  • (4)
    $R$ is called almost Gorenstein ([goto2015almost, Section 10]) if there exists an $R$-monomorphism $\phi:R\hookrightarrow\omega_{R}(-{a_{R}})$ of degree 0 such that $C:=\operatorname{coker}(\phi)$ is either the zero module or an Ulrich $R$-module, i.e. $C$ is Cohen–Macaulay and $\mu_{R}(C)=e(C)$. Here, $\mu_{R}(C)$ (resp. $e(C)$) denotes the number of elements in a minimal generating system of $C$ (resp. the multiplicity of $C$ with respect to ${\mathfrak{M}}_{R}$).
  • (5)
    $R$ is generically Gorenstein if $R_{\mathfrak{p}}$ is Gorenstein for every ${\mathfrak{p}}\in{\text{Ass}}(R)$. Notice that every domain is generically Gorenstein. When $R$ is generically Gorenstein, it is known that $\omega_{R}$ can be identified with a graded ideal in $R$ which contains a non-zero divisor of $R$ (see [bruns1998cohen, Proposition 3.3.18 (b)]).
In this paper, the following result is used to compare the relative strengths of the nearly Gorenstein and almost Gorenstein conditions.
Theorem 2.1 ([higashitani2016almost, Theorem 4.7]).
Let $R$ be a Cohen–Macaulay standard graded domain over an algebraically closed field $R_{0}=\mathbb{k}$ with characteristic 0. If $R$ is almost Gorenstein with $s(R)\geq 2$, then $R$ is pseudo-Gorenstein.
3. The relationship between pseudo-Gorensteinness and the canonical trace
In this section, we first prove Theorem 3.1, which can be applied to a general $R$-module $M$. In particular, by applying this proposition to the case where $M$ is the canonical module $\omega_{R}$, we prove Theorem A. Following the previous section, let $\mathbb{k}$ be a field.
[…]
Let $R$ be a graded ring. For a non-zero $\mathbb{Z}$-graded $R$-module $M$, we express as $\text{indeg}(M)<\infty$ if $\operatorname{min}\{i\in\mathbb{Z}:M_{i}\neq 0\}\in\mathbb{Z}$, and we set $\text{indeg}(M):=\operatorname{min}\{i\in\mathbb{Z}:M_{i}\neq 0\}$.
The following fundamental lemma plays an important role in proving our main result.
[…]
Let $R$ be a graded ring and let $I\subseteq R$ be a graded ideal of $R$ such that $\text{indeg}(I)<\infty$ and there exists homogeneous non-zero divisors $f_{1},f_{2}\in I$ with $f_{2}\notin Rf_{1}$. Let $M$ be a non-zero $\mathbb{Z}$-graded $R$-module and let $\theta_{1},\theta_{2}\in R_{\text{indeg}({\mathfrak{M}}_{R})}$ be an $R$-sequence. Suppose that $M$ contains, as a direct summand, a graded $R$-submodule that isomorphic to $I$. Then we have $\theta_{i}\frac{f_{2}}{f_{1}}\notin R$ for some $i=1,2$.
Proof.
Suppose that $\theta_{i}\frac{f_{2}}{f_{1}}\in R$ for each $i=1,2$. Then there are $r_{1},r_{2}\in R$ such that $\theta_{1}f_{2}=f_{1}r_{1}$ and $\theta_{2}f_{2}=f_{1}r_{2}$. Thus, since $\theta_{2}(f_{1}r_{1})=(f_{2}\theta_{1})\theta_{2}=(f_{2}\theta_{2})\theta_{1}=(f_{1}r_{2})\theta_{1}$, we have $f_{1}r_{1}\theta_{2}=f_{1}r_{2}\theta_{1}$. Since $f_{1}\neq 0$ is a non-zero divisor of $R$, we have $r_{1}\theta_{2}=r_{2}\theta_{1}$. Moreover, we obtain $r_{1}=r\theta_{1}$ for some $r\in R$ because $\theta_{2}+(\theta_{1})\in R/(\theta_{1})$ is a non-zero divisor of $R/(\theta_{1})$. Substituting this into $f_{2}\theta_{1}=f_{1}r_{1}$, we get $f_{2}\theta_{1}=rf_{1}\theta_{1}$. Since $\theta_{1}$ is a non-zero divisor of $R$, we have $f_{2}=rf_{1}$. This contradicts $f_{2}\notin Rf_{1}$. ∎
[…]
Let $R$ be a graded ring and let $M$ be a $\mathbb{Z}$-graded $R$-module. For $i\in\mathbb{Z}$, we say that $M$ has an $R(i)$-free summand if there exists an $R$-epimorphism $\phi\in\text{Hom}_{R}^{i}(M,R)$.
We establish the following, which presents Theorem A in a more general form.
Theorem 3.1.
Let $R$ be a graded ring and let $I\subseteq R$ be a graded ideal of $R$ such that $\text{indeg}(I)<\infty$ and there exists homogeneous $R$-regular element $f\in I_{\text{indeg}(I)}$ such that $Rf\neq I$. Let $M$ be a non-zero $\mathbb{Z}$-graded $R$-module with $M_{\text{indeg}(M)}\cong I_{\text{indeg}(I)}$. Suppose that $M$ contains, as a direct summand, a graded $R$-submodule that isomorphic to $I$ such that $[\textrm{tr}_{R}(M/I)]_{\text{indeg}({\mathfrak{M}}_{R})}=0$. If $\textrm{tr}_{R}(M)\supseteq(\theta_{1},\theta_{2})R$ for some $R$-sequence $\theta_{1},\theta_{2}\in R_{\text{indeg}({\mathfrak{M}}_{R})}$ and $\dim_{\mathbb{k}}(M_{\text{indeg}(M)})=1$, then $M$ has an $R(-\text{indeg}(M))$-free summand.
Proof.
Put $a=\text{indeg}({\mathfrak{M}}_{R})$ and $b=\text{indeg}(M)$. Take a graded ideal $I$ of $R$ and a homogeneous non-zero divisor $f=f_{1}\in I_{\text{indeg}(I)}$ that satisfy the assumptions of the statement. Since $Rf_{1}\neq I$, we can choose a homogeneous element $f_{2}\in I\setminus Rf_{1}$. We can take a homogeneous $R$-generating system $G(I)$ of $I$ such that $G(I)\setminus\{f_{1},f_{2}\}\subseteq I\setminus(Rf_{1}+Rf_{2})$. Then we have $\theta_{i}\frac{f_{2}}{f_{1}}\notin R$ for some $1\leq i\leq 2$ by Section missing 3.
Note that there exists a graded $R$-submodule $N$ of $M$ such that $M\cong I\oplus N$ as an $R$-module. Then, notice that $M_{b}=I_{b}\cong\mathbb{k}f_{1}$ as $\mathbb{k}$-vector space and $\operatorname{deg}f_{1}<\operatorname{deg}g$ for any homogeneous element $g\in(G(I)\setminus\{f_{1}\})\cup N$. Moreover, since $\operatorname{Hom}_{R}^{*}(I\oplus N,R)\cong\operatorname{Hom}_{R}^{*}(I,R)\oplus\operatorname{Hom}_{R}^{*}(N,R)$, it follows that $\textrm{tr}_{R}(M)=\textrm{tr}_{R}(I)+\textrm{tr}_{R}(N)$. Therefore, from $\theta_{i}\in\textrm{tr}_{R}(M)$, there exist $\phi_{1}\in\operatorname{Hom}_{R}^{\text{indeg}({\mathfrak{M}}_{R})-\operatorname{deg}(f_{1})}(I,R)$ and two finite subsets $A\subseteq\operatorname{Hom}_{R}^{*}(I,R)$ and $B\subseteq\operatorname{Hom}_{R}^{*}(X,R)$ such that $\theta_{i}\in\phi_{1}(f_{1})+\sum_{\phi\in A}\phi(G(I)\setminus\{f_{1}\})+\sum_{\phi\in B}\phi(N)$. Here, we have $\sum_{\phi\in B}\phi(N)=0$. Indeed, if $\sum_{\phi\in B}\phi(N)\neq 0$, then there exists some $\phi\in B$ and $n\in N$ such that $0\neq\phi(n)\in\mathbb{k}\theta_{i}\cong R_{\text{indeg}({\mathfrak{M}}_{R})}$. This leads to a contradiction because $00\neq[\textrm{tr}_{R}(N)]_{\text{indeg}({\mathfrak{M}}_{R})}=[\textrm{tr}_{R}(M/I)]_{\text{indeg}({\mathfrak{M}}_{R})}=$ 0. Thus, we obtain the following result.
(1) $\theta_{i}\in\phi_{1}(f_{1})+\sum_{\phi\in A}\phi(G(I)\setminus\{f_{1}\}).$
Here, suppose that $\sum_{\phi\in A}\phi(G(I)\setminus\{f_{1}\})=0$. In this case, since $\theta_{i}=\phi_{1}(f_{1})$ and $f_{1},f_{2},f_{1}f_{2}\in I\subseteq M$, we have $\theta_{i}\frac{f_{2}}{f_{1}}=\frac{\phi_{1}(f_{1})f_{2}}{f_{1}}=\frac{\phi_{1}(f_{1}f_{2})}{f_{1}}=\frac{\phi_{1}(f_{2})f_{1}}{f_{1}}=\phi_{1}(f_{2})\in R$. This contradicts to $\theta_{i}\frac{f_{2}}{f_{1}}\notin R$.
Thus we may assume that $\sum_{\phi\in A}\phi(G(I)\setminus\{f_{1}\})\neq 0$. Then, there exists a $\phi\in\text{Hom}^{*}_{R}(I,R)$ and a homogeneous element $f\in G(I)\setminus\{f_{1}\}$ of $M$ such that $\phi(f)$ is a non-zero homogeneous element of $R$. Since (missing) 1, we have $0\neq\phi(f)\in R_{a}$. Moreover, we can check $\phi(f_{1})\neq 0$. Indeed, if $\phi(f_{1})=0$, Noting the isomorphism $\text{Hom}_{R}(Rf_{1},R)\cong(Rf_{1})^{-1}$ and $\phi\neq 0$, there exists some $0\neq g\in(Rf_{1})^{-1}\subseteq Q(R)$ such that $gf_{1}=\phi|_{Rf_{1}}(f_{1})=\phi(f_{1})=0$. Since $f_{1}$ is a non-zero divisor of $R$, we have $g=0$, yields a contradiction.
Thus, $\phi(f_{1})\neq 0$. Consequently, $0\neq\phi(f)\in R_{a}$, and since $\operatorname{deg}(f_{1})<\operatorname{deg}(f)$, it follows that $\operatorname{deg}(\phi(f_{1}))<\operatorname{deg}(\phi(f))=a$. Then, since $\phi(f_{1})\in\mathbb{k}\setminus\{0\}$, we have $\phi\in\text{Hom}_{R}^{-\text{indeg}(M)}(M,R)$ and $\phi$ is epimorphism. Therefore, $M$ has an $R(-\text{indeg}(M))$-free summand. ∎
[…]
The above result is a substantial generalization of [miyashita2024comparing, Theorem 3.5]. In fact, from [miyashita2024linear, Proposition 4.4], we see that [miyashita2024comparing, Theorem 3.5] states the following: if $R$ is a semi-standard graded ring and $M=I$ is a non-principal ideal of $R$, then under the conditions $0pt(R)\geq 2$ and $R$ satisfies ()($\natural$)( ♮ ), we have $\dim_{\mathbb{k}}\bigl{(}I_{\text{indeg}(I)}\bigr{)}\geq 2$. Since $0pt(R)\geq 2$ and $R$ satisfies ()($\natural$)( ♮ ) imply $\text{grade}\bigl{(}R[\omega_{R}]_{\text{indeg}({\mathfrak{M}}_{R})}\bigr{)}\geq 2$, it follows that [miyashita2024comparing, Theorem 3.5] can be deduced from Theorem 3.1.
The following statement, which is a non-Cohen–Macaulay version of Theorem A and is also applicable to rings that are not generically Gorenstein, is obtained as a consequence of Theorem 3.1. For a Noetherian graded ring $R$ admits a graded canonical module $\omega_{R}$, $R$ is called quasi-Gorenstein if $\omega_{R}\cong R$ as a graded $R$-module. Note that $R$ is quasi-Gorenstein if and only if $\textrm{tr}(\omega_{R})=R$ ([aoyama1985endomorphism, Corollary 3.5]).
Theorem 3.2.
Let $R$ be a Noetherian graded ring. Suppose that $\omega_{R}$ admits a graded ideal $I\subseteq R$ as a direct summand, where $I_{\text{indeg}(I)}\cong[\omega_{R}]_{-a_{R}}$, $[\textrm{tr}_{R}(\omega_{R}/I)]_{\text{indeg}({\mathfrak{M}}_{R})}=0$ and $\text{grade}(RI_{\text{indeg}(I)})>0$. If $\text{grade}(R[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}({\mathfrak{M}}_{R})})\geq 2$ and $\dim_{\mathbb{k}}({[\omega_{R}]}_{-a_{R}})=1$, then $R$ is quasi-Gorenstein.
Proof.
We take a finitely generated graded ideal $I$, which is a direct summand of $\omega_{R}$ and satisfies the assumptions of the statement. Then, there exists an $R$-submodule $N$ of $\omega_{R}$ such that there is an isomorphism of graded modules $\omega_{R}\cong I\oplus N$. Since $\text{grade}(RI_{\text{indeg}(I)})>0$, there exists a non-zero divisor $f\in I_{\text{indeg}(I)}$. Suppose that $I\neq Rf$. Then, by Theorem 3.1, $\omega_{R}$ has an $R(a_{R})$-free summand. Consequently, we obtain $\textrm{tr}_{R}(\omega_{R})=\textrm{tr}_{R}(R(a_{R}))=R$. By [aoyama1985endomorphism, Proposition 3.3], $R$ is quasi-Gorenstein, i.e., $R\cong\omega_{R}$. In this case, the $R$-module isomorphism $I\oplus N\cong R$ implies that it is generated by a single element as an $R$-module. However, since $I$ has at least two minimal generators, including $f$, this leads to a contradiction (see [matsumura1989commutative, Theorem 2.3]). Therefore, we obtain $I=Rf\cong R$. Since $\textrm{tr}_{R}(\omega_{R})=\textrm{tr}_{R}(R)+\textrm{tr}_{R}(N)=R$, by applying [aoyama1985endomorphism, Proposition 3.3] again, we conclude that $R$ is quasi-Gorenstein, which means that $R\cong\omega_{R}$. ∎
As a special case of Theorem 3.2, we obtain the various corollaries described below. First, by imposing the condition that $R$ is generically Gorenstein in Theorem 3.2, we obtain the following result. Here, a Noetherian graded ring $R$ is said to be generically Gorenstein if $R_{\mathfrak{p}}$ is Gorenstein for every ${\mathfrak{p}}\in{\text{Ass}}(R)$.
[…]
Let $R$ be a Noetherian generically Gorenstein graded ring with $\text{grade}(R[\omega_{R}]_{-a_{R}})>0$ and $\text{grade}(R[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}({\mathfrak{M}}_{R})})\geq 2$. If $\dim_{\mathbb{k}}([\omega_{R}]_{-a_{R}})=1$, then $R$ is quasi-Gorenstein.
Proof.
Since $R$ is generically Gorenstein, there exists a graded ideal $I$ of $R$ such that $\omega_{R}\cong I$ (see [endo2023rings, Proposition 2.3]). In particular, $\omega_{R}$ has $I$ as a direct summand, and $\omega_{R}/I=0$. Hence, by applying Theorem 3.2, we see that $R$ is quasi-Gorenstein. ∎
Furthermore, by assuming the Cohen–Macaulay condition in Theorem 3.2 and Section missing 3 respectively, we obtain the following two results.
[…]
Let $R$ be a Cohen–Macaulay graded ring. Suppose that $\omega_{R}$ admits a graded ideal $I\subseteq R$ as a direct summand, where $I_{\text{indeg}(I)}\cong[\omega_{R}]_{-a_{R}}$, $[\textrm{tr}_{R}(\omega_{R}/I)]_{\text{indeg}({\mathfrak{M}}_{R})}=0$ and $\text{grade}(RI_{\text{indeg}(I)})>0$. If $\text{grade}(R[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}({\mathfrak{M}}_{R})})\geq 2$ and $R$ is pseudo-Gorenstein, then $R$ is Gorenstein.
[…]
Let $R$ be a Cohen–Macaulay generically Gorenstein graded ring. Assume
$\text{grade}(R[\omega_{R}]_{-a_{R}})>0,\;\text{grade}(R[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}({\mathfrak{M}}_{R})})\geq 2.$
Then $R$ is pseudo-Gorenstein if and only if $R$ is Gorenstein.
[…]
There are infinitely many counterexamples if we remove the assumption
$\text{grade}\bigl{(}R[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}({\mathfrak{M}}_{R})}\bigr{)}\geq 2$
from Section missing 3. Indeed, any 1-dimensional pseudo-Gorenstein Cohen–Macaulay graded domain (e.g., numerical semigroup ring) $R$ satisfies all the assumptions except $\text{grade}\bigl{(}R[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}({\mathfrak{M}}_{R})}\bigr{)}\geq 2$. Hence, if $R$ is not Gorenstein, then $R$ is a counterexample.
[…]
Let $R$ be a Cohen–Macaulay generically Gorenstein graded ring with $\text{grade}(R[\omega_{R}]_{-a_{R}})>0$  (e.g., $R$ is a domain), $0pt(\mathbb{k}[R_{\text{indeg}({\mathfrak{M}}_{R})}])\geq 2$ (e.g., $R$ is a semi-standard graded ring with $\dim(R)\geq 2$) and $\sqrt{R[\textrm{tr}_{R}(\omega_{R})]_{{\text{indeg}({\mathfrak{M}}_{R})}}}\supseteq{\mathfrak{M}}_{R}$ (e.g., $R$ is nearly Gorenstein with $0pt(\mathbb{k}[R_{\text{indeg}({\mathfrak{M}}_{R})}])\geq 2$). Then
  • (1)
    $R$ is pseudo-Gorenstein if and only if $R$ is Gorenstein;
  • (2)
    If the Cohen–Macaulay type $r(R)$ of $R$ is 2, then $R$ is level.
Proof.
We show (1). Suppose that $R$ is pseudo-Gorenstein. We may assume that $R_{0}=\mathbb{k}$ is infinite (see [miyashita2024linear, Remark 2.4 (4)]). Note that $\sqrt{R[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}({\mathfrak{M}}_{R})}}$ is either $R$ or ${\mathfrak{M}}_{R}$. In the former case, we have $\textrm{tr}_{R}(\omega_{R})=R$, which implies that $R$ is Gorenstein. In the latter case, since $0pt\bigl{(}\mathbb{k}[R_{\text{indeg}({\mathfrak{M}}_{R})}]\bigr{)}\geq 2$ and $\mathbb{k}$ is an infinite field, we can choose an $R$-regular sequence of length 2 in $[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}({\mathfrak{M}}_{R})}$ by the same argument used in the proof of [bruns1998cohen, Proposition 1.5.12]. Thus, $\text{grade}\bigl{(}R[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}({\mathfrak{M}}_{R})}\bigr{)}\geq 2$. Applying Section missing 3 in this situation completes the proof. We show (2). Notice that $r(R)$ is equal to the number of elements in a minimal generating system of $\omega_{R}$. If $R$ is not level of type 2, $R$ is pseudo-Gorenstein, and it follows that $R$ is Gorenstein by Section missing 3 (2). Hence, in particular, $R$ is level, a contradiction. ∎
By applying this to a graded domain, we obtain the following corollaries.
[…]
The following holds:
  • (1)
    Every nearly Gorenstein and pseudo-Gorenstein graded domain $R$ with $0pt(\mathbb{k}[R_{\text{indeg}({\mathfrak{M}}_{R})}])\geq 2$ is Gorenstein;
  • (2)
    Every pseudo-Gorenstein semi-standard graded domain with ()($\natural$)( ♮ ) of dimension $\geq 2$ is Gorenstein.
4. Some applications
In this section, we present several applications of the results obtained in the previous section. First, we prove the following statement, which generalizes [miyashita2024comparing, Corollary 3.7 and Theorem 6.1].
[…]
The following holds:
  • (1)
    All type 2 semi-standard graded Cohen–Macaulay domains satisfy ()($\natural$)( ♮ ) of dimension $\geq$ 2 are level;
  • (2)
    All type 2 nearly Gorenstein semi-standard graded domains of dimension $\geq$ 2 are level;
  • (3)
    Let $R$ be a Cohen–Macaulay standard graded domain with $\dim(R),s(R)\geq 2$. Assume that $R_{0}=\mathbb{k}$ is algebraically closed field with $\operatorname{char}(\mathbb{k})=0$. Then the following conditions are equivalent:
    • (a)
      $R$ satisfies ()($\natural$)( ♮ ) and almost Gorenstein;
    • (b)
      $R$ is nearly Gorenstein and almost Gorenstein;
    • (c)
      $R$ is Gorenstein.
  • (4)
    All nearly Gorenstein and almost Gorenstein standard graded domains over an algebraically closed field $\mathbb{k}$ with $\operatorname{char}(\mathbb{k})=0$ of dimension $\geq 2$ are level.
Proof.
(1) follows from Section missing 3 (2) and Section missing 3 (2). (2) is a special case of (1). We show (3). Since $(c)\Rightarrow(b)\Rightarrow(a)$ is clear, we only need to show $(a)\Rightarrow(c)$. By [higashitani2016almost, Theorem 4.7], $R$ is pseudo-Gorenstein, and hence $R$ is Gorenstein by Section missing 3 (2). We show (4). When $s(R)=1$ (i.e., $R$ has minimal multiplicity), it is known that $R$ is level. Thus it follows from (3) that $R$ is Gorenstein. ∎
Finally, we discuss an application to Veronese subalgebras. We begin by briefly recalling the definition.
[…]
Let $k\in\mathbb{Z}_{>0}$. For a graded ring $R=\bigoplus_{i\geq 0}R_{i}$, the $k$-th Veronese subalgebra of $R$ is defined as $R^{(k)}=\bigoplus_{i\geq 0}R_{ik}$. Similarly, for a graded $R$-module $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$, we define $M^{(k)}:=\bigoplus_{i\in\mathbb{Z}}M_{ik}$.
For a Noetherian graded ring $R$, let ${\text{Assh}}(R)$ denote
${\text{Assh}}(R)=\{{\mathfrak{p}}\in{\text{Ass}}(R):{\mathfrak{p}}\text{ is a graded ideal of $R$ with }\dim(R)=\dim(R/{\mathfrak{p}})\}.$
The following is the non-Cohen–Macaulay version of [miyashita2024linear, Proposition 5.5].
[…]
Let $R$ be a Noetherian semi-standard graded ring such that $00pt(R)>$ 0 and $R_{\mathfrak{p}}$ is Gorenstein for any ${\mathfrak{p}}\in{\text{Assh}}(R)$. Let $I$ be a graded ideal generated by a subset of $R_{1}$. Put ${\mathfrak{N}}_{R}=[{\mathfrak{M}}_{R}]_{1}R$. If $\textrm{tr}_{R}(\omega_{R})\supseteq I$, then we have $\textrm{tr}_{R^{(k)}}(\omega_{R^{(k)}})\supseteq({\mathfrak{N}}_{R}^{k-1}I)^{(k)}$ for any $k>0$.
Proof.
By [endo2023rings, Proposition 2.3], there exists a graded ideal $I$ of $R$ such that $\omega_{R}\cong I$. Since $R$ is semi-standard graded and $00pt(R)>$ 0, we can take a non-zero-divisor $z\in R_{1}$ (see [miyashita2024linear, Lemma 5.2]). Considering $z^{-a_{R}-\text{indeg}(I)}\in Q(R)$, we have $I\cong z^{-a_{R}-\text{indeg}(I)}I$. Thus, there exists a graded fractional ideal $J:=z^{-a_{R}-\text{indeg}(I)}I$ of $R$ such that $\omega_{R}\cong J$ and $-a_{R}=\text{indeg}(J)$. Since $\omega_{R^{(k)}}\cong(\omega_{R})^{(k)}$ (see [goto1978graded, Corollary (3.1.3)]) for any $k>0$, we have $\textrm{tr}_{R^{(k)}}(\omega_{R^{(k)}})=\textrm{tr}_{R^{(k)}}(J^{(k)})\supseteq({\mathfrak{N}}_{R}^{k-1}I)^{(k)}$ by [miyashita2024linear, Proposition 5.3]. ∎
With the above preparations, we can now prove the following.
Theorem 4.1.
Let $R$ be a Noetherian semi-standard graded generically Gorenstein ring such that $[\omega_{R}]_{-a_{R}}$ contains an $R$-regular element. Assume that $a_{R}\in k\mathbb{Z}$ and $0pt(R^{(k)})\geq 2$ for some $k\in\mathbb{Z}_{>0}$. If $\dim_{\mathbb{k}}([\omega_{R}]_{-a_{R}})=1$ and there exists a system of parameter $\theta_{1},\theta_{2}\in R_{1}$ such that $\textrm{tr}_{R}(\omega_{R})\supseteq(\theta_{1},\theta_{2})R$, then $R^{(k)}$ is quasi-Gorenstein.
Proof.
Put ${\mathfrak{N}}_{R}=(R_{1})R$. Then we have $\textrm{tr}_{R^{(k)}}(\omega_{R^{(k)}})\supseteq({\mathfrak{N}}_{R}^{k-1}(\theta_{1},\theta_{2})R)^{(k)}=({\mathfrak{N}}_{R}^{k-1}(\theta_{1},\theta_{2})R)\cap R^{(k)}$ by Section missing 4. We have $\sqrt{(\mathfrak{n}_{R}^{k-1}(\theta_{1},\theta_{2})R)\cap R^{(k)}}=\sqrt{(\mathfrak{n}_{R}^{k-1}(\theta_{1},\theta_{2})R)}\cap R^{(k)}=\mathfrak{m}_{R}\cap R^{(k)}=\mathfrak{m}_{R^{(k)}}$, which implies that $(\mathfrak{n}_{R}^{k-1}(\theta_{1},\theta_{2})R)\cap R^{(k)}$ is an $\mathfrak{m}_{R^{(k)}}$-primary ideal of $R^{(k)}$ generated by the subset $\{f^{k-1}\theta_{i}\in R^{(k)}:f\in R_{1},\ i=1,2\}$ of $[R^{(k)}]_{1}=R_{k}$. Thus, $R^{(k)}$ satisfies property ()($\natural$)( ♮ ). Since $R$ is a generically Gorenstein semi-standard graded ring, so is $R^{(k)}$ (noting that $[R^{(k)}]_{1}=R_{k}$). Moreover, from $\omega_{R^{(k)}}=(\omega_{R})^{(k)}$ and the assumption $a_{R}\in k\mathbb{Z}$, it follows that $a_{R^{(k)}}=a_{R}$. Consequently, $\dim_{\mathbb{k}}([\omega_{R^{(k)}}]_{-a_{R^{(k)}}})=1$. Since $R^{(k)}$ satisfies ()($\natural$)( ♮ ), it follows that $R^{(k)}\otimes_{\mathbb{k}}\mathbb{k}(x)$ also satisfies ()($\natural$)( ♮ ) (see [miyashita2024linear, Remark 3.2 (3)]). Noting that $\omega_{R\otimes_{\mathbb{k}}\mathbb{k}(x)}=\omega_{R}\otimes_{\mathbb{k}}\mathbb{k}(x)$ (see [aoyama1985endomorphism, Theorem 4.1]), we may assume, by replacing $R$ with $R\otimes_{\mathbb{k}}\mathbb{k}(x)$ if necessary, that $\mathbb{k}$ is an infinite field. Therefore, since $R^{(k)}$ satisfies ()($\natural$)( ♮ ) and $0pt(R^{(k)})\geq 2$, we obtain $\text{grade}(R[\textrm{tr}_{R}(\omega_{R})]_{\text{indeg}(\mathfrak{m}_{R})})=\text{grade}(R[\textrm{tr}_{R}(\omega_{R})]_{1})\geq 2$. Since $R^{(k)}$ is generically Gorenstein (as is $R$), applying Section missing 3 to $R^{(k)}$ shows that $R^{(k)}$ is quasi-Gorenstein. ∎
[…]
Let $R$ be a Noetherian semi-standard graded generically Gorenstein ring such that $[\omega_{R}]_{-a_{R}}$ contains an $R$-regular element. Suppose that $a_{R}\in k\mathbb{Z}$ and $0pt(R^{(k)})\geq 2$ for some $k\in\mathbb{Z}_{>0}$. Additionally, assume that $\sqrt{[\textrm{tr}_{R}(\omega_{R})]_{1}R}\supseteq{\mathfrak{M}}_{R}$ (e.g., $\textrm{tr}_{R}(\omega_{R})\supseteq{\mathfrak{M}}_{R}$) and that $\dim_{\mathbb{k}}([\omega_{R}]_{-a_{R}})=1$. Then $R^{(k)}$ is quasi-Gorenstein. In particular, if $R^{(k)}$ is Cohen–Macaulay, then it is Gorenstein.
Acknowledgement
I would like to thank Kazufumi Eto, Akihiro Higashitani, Shinya Kumashiro, Naoyuki Matsuoka, Takahiro Numata, Kei-ichi Watanabe and Kohji Yanagawa for their useful comments. I would also like to express my gratitude to Max Kölbl for his assistance in proofreading the English manuscript. This work was supported by JST SPRING, Japan Grant Number JPMJSP2138.